OpenStudy (anonymous):

write an equation for the line passing through B(0,3) and C(-3,4)

OpenStudy (texaschic101):

first we need to find the slope by using the slope formula. slope(m) = (y2 - y1) / (x2 - x1) (0,3)...x1 = 0 and y1 = 3 (-3,4)...x2 = -3 and y2 = 4 now we sub slope(m) = (4 - 3) / (-3 - 0) slope(m) = 1/-3 or -1/3 now we use y = mx + b...we know the slope, but we have to find b (the y intercept) slope(m) = -1/3 choose either of your points...I choose (0,4)...x = 0 and y = 4 now we sub 4 = -1/3(0) + b 4 = b so our equation is : y = -1/3x + 4

OpenStudy (texaschic101):

do you have any questions at all because I will be happy to explain anything I did

OpenStudy (anonymous):

You are making this so much easier to understand. haha thank you :)